OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (5, 1) and (–4, 7). Answers: 1. Y= - 2/3x + 13/3 2. Y= -2/3x - 13/3 3. Y= 2/3x - 13/3 4. Y= 2/3x + 13/3

OpenStudy (hockeychick23):

y= -2/3x+13/3

OpenStudy (anonymous):

So A? (: Thank you. Do you have time for one more equation thing?

OpenStudy (shamim):

Y-y1=(y1-y2)/(x1-x2)(x-x1)